OpenStudy (anonymous):

Find the reference angle of 1020 degrees

OpenStudy (anonymous):

1020 - 360 = 660 660 - 360 = 300 360 - 300 = 60 so it is 60 degrees

OpenStudy (anonymous):

oh,, thanks, xD

OpenStudy (anonymous):

wait,, how'd that happen?

OpenStudy (anonymous):

@RED123url ??

OpenStudy (anonymous):